OpenStudy (anonymous):

Factor completely 121x2 – 88x + 16

OpenStudy (anonymous):

this one you recognize as a "perfect square" because \[121x^2=(11x)^2, 16=(-4)^2, 2\times 11\times 4=88\] so you get \[(11x-4)^2\]